# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

### v1

@@ 5.2.1. Basic properties
Recall that a bielliptic surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=1$ and $h^{2}({\mathcal{O}}_{S})=0$. In particular, the exponential sequence reads
@@
where $\operatorname{Pic}^{0}(S)=H^{1}({\mathcal{O}}_{S})/H^{1}(S,\mathbb{Z})$ is the group of algebraically trivial divisors modulo linear equivalence and has the structure of an elliptic curve. There exist two elliptic curves $A$, $B$, and an abelian group $G$ acting on $A$ and $B$, such that ([31, Proposition 1.1])
@@
    $A/G$ is elliptic and $B/G\cong\mathbb{P}^{1}$;
@@
    $S\cong(A\times B)/G$, where $G$ acts on $A\times B$ componentwise;
see [1]. We follow the convention in [31] by abuse of notation to denote by $A$ (resp. $B$) the class (in $\operatorname{Num}(S)$, $H^{2}(S,\mathbb{Z})$, or $H^{2}(S,{\mathbb{Q}})$) of a smooth fiber of the natural projection $\Psi:S\to B/G$ (resp. $\Phi:S\to A/G$). These classes satisfy
@@
where $\gamma$ is the order of $G$. Let $\lambda A$, $\mu B$ be a basis for $\operatorname{Num}(S)$ ($\lambda,\mu\in{\mathbb{Q}}$); see [31, Theorem 1.4] for the values of $\lambda$ and $\mu$.
We need the following lemma to compute the dimension of linear systems.
Lemma 5.10 ([31, Lemma 1.3]).
Let $D$ be a divisor of numerical class $sA+tB$ with $s,t\in{\mathbb{Q}}$ on a bielliptic surface. The following properties hold:
@@
The numerical class of $\beta$ in $\operatorname{Num}(S)$ is of the form $a\lambda A+b\mu B$ with $a,b\in\mathbb{Z}_{>0}$. By Lemma 5.10 (i) and (iii),
@@
Any non-integral curve in $|d\beta|$ is in the image of a morphism from a scheme $P$ to $|d\beta|$, where $P$ parametrizes the triples $(C,D,N)$ satisfying $N\in\operatorname{Pic}^{0}(S)$, $C\in|C_{1}+N|$, and $D\in|C_{2}-N|$ for two nonzero effective divisors $C_{1}$, $C_{2}$ such that $C_{1}+C_{2}=d\beta$. By Lemma 4.5 (or Lemma 5.10), the images of finitely many such morphisms cover $|d\beta|\setminus|d\beta|^{\mathrm{int}}$. Thus, it suffices to estimate all possible values of
@@
Therefore, by the proof of [29, (13)] (or Case 1.1 in §5.1.2), we have
(5.28) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1=C_{1}\cdot C_{2}\geq d\sqrt{\beta^{2}}-1.$
@@
If $b_{2}>0$, then it follows from $b_{1}>0$ and (5.29) that
(5.30) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $=C_{1}\cdot C_{2}-\dim|C_{1}|$
$\geq dab_{1}\lambda\mu\gamma-(b_{1}\mu\gamma-1)$
$\geq(da\lambda-1)\mu\gamma+1\quad(\text{if }da\lambda>1).$
If $b_{2}=0$, then

### v2

@@ 5.2.1. Basic properties
Recall that a bielliptic surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=1$ and $h^{2}({\mathcal{O}}_{S})=0$. In this case, the exponential sequence reads
@@
where $\operatorname{Pic}^{0}(S)=H^{1}(S,{\mathcal{O}}_{S})/H^{1}(S,\mathbb{Z})$ is the group of algebraically trivial divisors modulo linear equivalence and has the structure of an elliptic curve. By [1, pp. 112–114] (see also [39, Proposition 1.1]), there exist two elliptic curves $A$, $B$, together with a finite Abelian group $G$ acting on $A$ and $B$, such that
@@
    $A/G$ is an elliptic curve and $B/G\cong\mathbb{P}^{1}$;
@@
    $S\cong(A\times B)/G$, where $G$ acts on $A\times B$ componentwise.
We follow the convention in [39] to denote also by $A$ (resp. $B$) the class in $\operatorname{Num}(S)$ of a smooth fiber of the natural projection $\Psi:S\to B/G$ (resp. $\Phi:S\to A/G$). Then
@@
where $\gamma$ is the order of $G$. There exist $\lambda,\mu\in{\mathbb{Q}}_{>0}$ such that the classes $\lambda A$ and $\mu B$ form a basis of $\operatorname{Num}(S)$ (see [39, Theorem 1.4] for the values of $\lambda$ and $\mu$).
We need the following lemma to compute the dimension of linear systems on $S$.
Lemma 5.6 ([39, Lemma 1.3]).
Let $D$ be a divisor of numerical class $uA+vB$ with $u,v\in{\mathbb{Q}}$ on a bielliptic surface $S$. The following properties hold:
@@
By Lemma 5.6 (i)–(iii), the numerical class of $\beta$ in $\operatorname{Num}(S)$ is of the form $a\lambda A+b\mu B$ with $a,b\in\mathbb{Z}_{>0}$, and
@@
Any non-integral curve in $|d\beta|$ is in the image of a morphism from a scheme $P$ to $|d\beta|$, where $P$ parametrizes the triples $(C,D,L)$ satisfying $L\in\operatorname{Pic}^{0}(S)$, $C\in|C_{1}+L|$, and $D\in|C_{2}-L|$ for two nonzero effective divisors $C_{1}$, $C_{2}$ such that $C_{1}+C_{2}=d\beta$. By Lemma 4.9 (or Lemma 5.6), the images of finitely many such morphisms cover $|d\beta|\setminus|d\beta|^{\mathrm{int}}$. Thus, it suffices to estimate all possible values of
@@
Therefore, by the proof of (5.18), we have
(5.25) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1=C_{1}\cdot C_{2}\geq d\sqrt{2\beta^{2}}-2.$
@@
If $b_{2}>0$, then it follows from Lemma 5.6 (i)–(iii), (5.26), and $b_{1}>0$ that
(5.27) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $=C_{1}\cdot C_{2}-\dim|C_{1}|-1$
$\geq dab_{1}\lambda\mu\gamma-(b_{1}\mu\gamma-1)-1$
$\geq(da\lambda-1)\mu\gamma\quad(\text{if }da\lambda\geq 1).$
If $b_{2}=0$, then as in (5.26), we have
@@
5.3. The case of K3 or Abelian surfaces
Now let $S$ be a K3 or an Abelian surface. The proof of Proposition 5.1 in this case is very similar to the proofs for Enriques and bielliptic surfaces (given in §5.1.2 and §5.2.2, respectively), so we keep it brief. Since the Kawamata–Viehweg vanishing theorem shows that Lemma 5.2 (i) is also valid for K3 and Abelian surfaces, we only need the following lemma (see [19, Theorem] for the K3 case, and [17, Proposition 2.3, Corollary 2.6] for the Abelian case).
Lemma 5.7.
Let $D$ be a nonzero effective nef divisor on a K3 or an Abelian surface $S$ with $D^{2}=0$. Then $D$ is numerically equivalent to $mE$ for some $m\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E$ with $E^{2}=0$. Moreover,
[…]
5.4. Some related stabilization questions
We collect some stabilization questions for moduli spaces of one-dimensional sheaves. Now $S$ is not necessarily a minimal surface of Kodaira dimension 0. For simplicity, we may assume $h^{1}({\mathcal{O}}_{S})=0$. We begin by providing a refined version of Conjecture 1.3, stated heuristically as follows.
Conjecture 5.8.
For fixed $i,j\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$, the perverse Hodge number $n_{h}^{i,j}$ of $h:M_{\beta,\chi}\rightarrow|\beta|$ stabilizes as $\beta$ becomes sufficiently positive. Furthermore, $n_{h}^{i,j}$ stabilizes to $H(q,t)]^{i,j}$ (defined as in Theorem 4.5) if $M_{\beta,\chi}$ is asymptotically irreducible.
In [36], the approach to Conjecture 1.3 for del Pezzo surfaces is to prove Conjecture 5.8. When $S$ is a K3 or an Abelian surface and $M_{\beta,\chi}$ is smooth, Shen–Yin [40] proved that the perverse Hodge numbers of $M_{\beta,\chi}$ are equal to its Hodge numbers using hyper-Kähler geometry and thus the stabilization of the perverse Hodge numbers follows from the closed formula of Hodge numbers of $S^{[n]}$ [12, Theorem 2.3.14]. Our Theorem 1.6 also shows the stabilization of perverse Hodge numbers when $S$ is an Enriques surface and $M_{\beta,\chi}$ is smooth.
Question 5.9.
For fixed $i,j\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$, does the intersection Hodge number $\dim_{\mathbb{Q}}I\!H^{i,j}(M_{\beta,\chi})$ stabilize as $\beta$ becomes sufficiently positive?
It seems that our method in this paper does not work for intersection Hodge numbers in the cases of Enriques and bielliptic surfaces. In [29], Mauri and Migliorini showed that Question 5.9 has a positive answer for K3 surfaces, based on the previously known case where $\chi=1$ and $S$ is a K3 surface.
The stabilization problems for moduli spaces of one-dimensional sheaves on $S$ can also be studied from a motivic perspective. Let $K_{0}(\mathrm{Var})$ be the Grothendieck ring of varieties, which is the free Abelian group generated by (the isomorphism classes of) varieties $[X]$ modulo the scissor relations
$[X]=[Y]+[Z],$
where $Y$ and $Z$ are disjoint locally closed subvarieties of $X$ with $X=Y\cup Z$. The product in $K_{0}(\mathrm{Var})$ is defined by taking the fiber product. Let $\mathbb{L}\in K_{0}(\mathrm{Var})$ be the class of the affine line $\mathbb{A}^{1}$. Denote by $\widehat{K}_{0}(\mathrm{Var})$ the completion of $K_{0}(\mathrm{Var})[\mathbb{L}^{-1}]$ with respect to the dimension filtration (see [8, p. 17] for more details).
Question 5.10.
For fixed $\chi\in\mathbb{Z}$, does the class $M_{\beta,\chi}]$ stabilize in $\widehat{K}_{0}(\mathrm{Var})$ as $\beta$ becomes sufficiently positive?
Coskun and Woolf [8] proved that when $S$ is a rational surface and $K_{S}\cdot H<0$, the class of the moduli space of $H$-stable torsion-free sheaves with fixed Chern character on $S$ stabilizes in $\widehat{K}_{0}(\mathrm{Var})$ as the determinant of sheaves tends to $\infty$. A positive answer to Question 5.10 would establish an analogous result in the rank 0 case.
A natural further question is whether there is a geometric interpretation for the stabilization of Betti numbers. One potential approach is to consider the tautological classes in $H^{*}(M_{\beta,\chi})$, i.e., the Künneth components of Chern classes of the universal sheaf (assuming its existence) on $M_{\beta,\chi}\times S$. Let $RH^{*}(M_{\beta,\chi})\subset H^{*}(M_{\beta,\chi})$ be the subring generated by the tautological classes. Although $RH^{*}(M_{\beta,\chi})\neq H^{*}(M_{\beta,\chi})$ in general (cf. [37, Theorem 5.10]), we expect the stabilization of Betti numbers of $M_{\beta,\chi}$ to be explained by the stabilization of $RH^{*}(M_{\beta,\chi})$. When $S$ is a del Pezzo surface and $M_{\beta,\chi}$ is smooth, one has $RH^{*}(M_{\beta,\chi})=H^{*}(M_{\beta,\chi})$ and the stable Betti numbers are determined by the combinatorics of tautological classes (see [36, Corollary 1.8]). If our expectation is true, the stabilization phenomenon in the del Pezzo case fits into a more general picture.
As for the intersection cohomology, the $k$-th stable intersection Betti number calculated in this paper also coincides with $b_{k}^{\infty}$. In what follows, we assume that $M_{\beta,\chi}$ is pure dimensional. Note that there is a homomorphism of graded vector spaces (e.g., [30, Exercise 6.7.1, Remark 6.7.2])
$\alpha:H^{*}(M_{\beta,\chi})\rightarrow I\!H^{*}(M_{\beta,\chi}).$
Thus we can define $RI\!H^{\ast}(M_{\beta,\chi})$ as the image of $RH^{\ast}(M_{\beta,\chi})$ under the map $\alpha$. Given our results, we also expect the stabilization of intersection Betti numbers to be explained by the stabilization of ${RI\!H}^{\ast}(M_{\beta,\chi})$, and therefore we ask the following question.
Question 5.11.
Is it true that $\dim_{\mathbb{Q}}{RI\!H}^{k}(M_{\beta,\chi})=b_{k}^{\infty}$ if $\beta$ is sufficiently positive?
Another direction is to explore the tautological relations in $RH^{\ast}(M_{\beta,\chi})$. When $S=\mathbb{P}^{2}$, the tautological relations are studied in [20], leading to the formulation of a strengthened version of the “P=C” conjecture proposed in [21]. For other surfaces, however, the problem remains unexplored.
More generally, one may consider similar questions for moduli spaces of one-dimensional sheaves on a projective surface with mild singularities. In this setting, we continue to use the notation $M_{\beta,\chi}$.
Question 5.12.
Fix $\chi\in\mathbb{Z}$ and let $\beta$ be a divisor on a simply connected projective surface with ADE singularities. For fixed $k\in\mathbb{Z}_{\geq 0}$, does the intersection Betti number $\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})$ stabilize as $\beta$ becomes sufficiently positive? Similarly, for fixed $i,j\in\mathbb{Z}_{\geq 0}$, does the intersection Hodge number $\dim_{\mathbb{Q}}I\!H^{i,j}(M_{\beta,\chi})$ stabilize?
Toward answering Question 5.12, we expect our strategy in this paper to work in some cases after suitable modifications, but there are some difficulties. For example, the relative Hilbert schemes of points over the linear system $|\beta|$ may not be equidimensional as the curves in $|\beta|$ may not be locally planar. Moreover, the Betti numbers (or intersection Betti numbers) of Hilbert schemes of points on a singular surface are still unknown.
Acknowledgments
The authors would like to thank Mirko Mauri, Weite Pi, and Junliang Shen for helpful discussions. FZ is grateful to her advisor Jun Li for his constant support, to János Kollár and Claire Voisin for inspiring discussions on the deformation theory of curve singularities, to Georg Oberdieck for suggesting considering the case of bielliptic surfaces, and to Songsong Huang for useful conversations.
FS is partially supported by the Fundamental Research Funds for the Central Universities, NSFC (No. 12025106) and Shaanxi NSF (No. 2025JC-QYCX-002). FZ is supported by the NSFC grants (No. 12121001 and No. 12425105) and LMNS.